The sidecap package (package option wide and the wide environment)

[figure]

Fig. 1: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

text area (odd page) margin area

[figure] Fig. 2: XYZ The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

Fig. 3: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

[figure]

margin area text area (even page)

[figure]

Fig. 4: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

Fig. 5: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345– 67890 jumps!

text area (odd page) margin area

[figure]